OpenStudy (anonymous):

Solve for z by using the quadratic formula: z^2 - 4z = 0. A. 0, -4 B. 0, 4 C. 1, 4 D. 1, -4

sam (.sam.):

z^(2)-4z=0 z(z-4)=0 z=0 z-4=0 z=4 z=0,4

hero (hero):

Yeah, exactly....forget the quad formula.

hero (hero):

It's clear that people would rather use methods other than the quad formula if necessary.

Directrix (directrix):

I was wondering if the work had to be shown with the answer. If that is the case, then the instructions "Solve for z by using the quadratic formula" would not have been followed. If only the answer has to be punched in somewhere, certainly factoring would be the efficient route.

hero (hero):

Using quad form should only be a last resort. When other methods have failed.